3. Determine the intensity of shear of an oil having viscosity =1 poise. The oil is used for lubricating the clearance between a shaft of diameter 10cm and its journal bearing. The clearance is 1.5 mm and the shaft rotates at 150 r.p.m


1
Expert's answer
2021-06-09T16:37:01-0400

Given:

"\mu=1\ poise=0.1N\cdot s/m^2"

Diameter of shaft "D=10cm=0.1m"

Distance between shaft and journal bearing "dy=1.5mm=1.5\cdot10^{-3}m"

Speed of shaft "N=150\ r.p.m."


Tangential speed of shaft is given by



"u=\dfrac{\pi DN}{60}=\dfrac{\pi\cdot0.1m\cdot150\ r.p.m.}{60}=\dfrac{\pi}{4}m/s"

Change of velocity between shaft and bearing



"du=u-0=\dfrac{\pi}{4}m/s"

Mathematically



"\tau=\mu\dfrac{du}{dy}=0.1N\cdot s/m^2\times\dfrac{\dfrac{\pi}{4}m/s}{1.5\cdot10^{-3}m}=""=\dfrac{50\pi}{3}N/m^2\approx52.360 N/m^2"